OpenStudy (anonymous):

Isaac is looking through binoculars on a whale watching trip when he notices a sea otter in the distance. If he is 20 feet above sea level in the boat, and the angle of depression is 30 degrees, how far away from the boat is the otter (to the nearest foot)?

OpenStudy (anonymous):

\[\tan 0=\frac{ 20 }{ x }\] so here we want x, therefore multiply both sides by x, to get \[x \tan 0=\frac{ 20 }{ x }*x\] \[x \tan 0=20\] divide both sides by tan o gives \[\frac{ x \tan 0 }{ \tan 0 }\] =\[\frac{ 20 }{ \tan 0 }\] w/c gives \[x=\frac{ 20 }{ \tan 0 }\] x=?

OpenStudy (anonymous):

did you get it? angle of depression \[0=30\] so x= 20/ tan 30 what did you get

OpenStudy (anonymous):

so is that the answer

OpenStudy (anonymous):

@heather040200

OpenStudy (anonymous):

you just have to do 20/tan 30 and youll get ur answer